Hello!
Find an equation of the line that satisfies the given conditions.
through (10, 2); perpendicular to the line x - 4y + 4 = 0
thanks!

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Step 1: Put your equation into slope-intercept form, that is, solve for y so that it looks like:

Then use the slope, m, of the given line that you can determine by inspection of the slope-intercept form of your equation and the fact that:



to determine the slope of the desired line.

Then use this new slope value, the given point, and the point-slope form of the equation of a line to derive the desired equation:



John


Answer by stanbon(75887)

Find an equation of the line that satisfies the given conditions.
through (10, 2); perpendicular to the line x - 4y + 4 = 0
-------------------------------------
1st: Find the slope of the given line.
x - 4y +4 = 0
4y = x+4
y = (1/4)x + 1
slope = 1/4
-------------------------
2nd: The slope of all lines that are perpendicular
to the given line is -4.
-------------------------
3rd: Find the equation of the perpendicular line:
Form y = mx+b
You know y = 2 when x = 10, and you know m = -4
Solve for "b".
2 = -4*10 + b
b = 42
---------------------
Equation:
y = -4x + 42
=====================
Cheers,
Stan H.